# Conformal Generative Modeling with Improved Sample Efficiency through Sequential Greedy Filtering

### v1

@@ 8 Conclusion
9 Reproducibility Statement
Our anonymized source code is available in the supplementary material. Detailed instructions for reproducing all experiments are provided in the README.md file at the top level of the code directory. All parameters for running the experiments can be found in the config folders within the script folders of each experiment folder. In addition, we will release our code on GitHub upon acceptance of our manuscript.
@@ Appendix A Relationship between Risk Control and Admissibility Control
Appendix B Bounding Conformal Admissibility by Learn-then-Test Admissibility
@@
We note that 17 is bounded from below by 00. We further bound 16 as follows
@@
where the last inequality holds by the learn-then-test guarantee 15. ∎
Appendix C Implementation Details
C.1 Choice of the Individual Admissibility Levels
@@
where $K$ is the amount of prediction steps (typically, $K=3$). We note that 18 is somewhat analagous to a Bonferroni correction in that it also equally distributes a correction. In practice, when working with a upper bound on the sample size max (outlined in § 4.2), however, lower admissibility levels can often be attained by different choice of $\alpha_{(s)}$. Specifically, we choose a larger risk level for the generation step ($s=0$) and distribute the remaining risk equally among the other step(s) ($s>0$). Specifically, we choose
@@
For all experiments (§ 5), we choose $M=5$.
C.2 Distance Upper Bound
@@ C.3 Grid Specification for CLM
Appendix D Experimental Details
D.1 Setup Details
@@
TriviaQA. We remove duplicates automatically rather than calibrating a parameter to do this, in order to improve statistical efficiency of both methods. For CLM, we set the diversity parameter $\lambda_{(2)}=0.5$ such that all duplicates are removed and for SCOPE-Gen, we use a fixed duplicate removal filter instead of a diversity filter at the end of the prediction pipeline.
@@ D.2 Examples
Appendix E Additional Experiments
E.1 Admissibility Analysis
We demonstrate the empirical admissibility levels for the MIMIC-CXR experiment at admissibility level $\alpha=0.3$, varying calibration set size $n$ and the used non-conformity measure. The rest of the experimental setup is identical to the setup used for all other experiments (see caption of Tab. 1)
As shown in the histograms of Fig. 3, for all non-conformity measures, the admissibility achieved by SCOPE-Gen (blue) is slightly conservative for low sample sizes and converges to the desired level as the calibration set size $n$ is increased from $300$ to $1200$. This is similar for CLM (orange), where the achieved admissibility (using the optimization strategy described in § 5) also becomes less conservative for large sample sizes. However, in contrast to SCOPE-Gen, CLM is generally more conservative. In addition, this conservativeness is highly influenced by the choice of non-conformity measure: While the CLM gap becomes tighter for the max and sum non-conformities, this is not the case for the count non-conformity, where a large gap persists even for $n=1200$.
@@ E.1 Admissibility Analysis
Figure 3: Admissibility Analysis for MIMIC-CXR. For all non-conformity scores, SCOPE-Gen (blue) becomes less conservative with respect to the desired admissibility level $1-\alpha$ (red dashed line) as the amount of calibration samples $n$ increases, as typical for methods based on conformal prediction. For CLM (orange), in contrast, conservativeness depends much on the chosen non-conformity measure.
E.2 Further Quantitative Evaluations
We demonstrate further experiments that showcase the performance for the different non-conformity measures (count and max ($\gamma=0.3$) in addition to sum ($\gamma=0.5$); see 10), including two more baselines (baselines 3) and 4) introduced in § 5).
We see in LABEL:tab:performance_further that typically either SCOPE-Gen or one of its ablations performs best in terms of the considered evaluation metrics. An exception occurs on the CNN/DM and Molecules experiment when the max non-conformity is used, where CLM performs best in terms of prediction set size (for CNN/DM, $2.751$ compared to $3.872$ achieved by “SCOPE-Gen gen. only”). Another noteworthy observation is that in some experiments, “SCOPE-Gen gen. only” generates the smallest prediction sets. For instance, on MIMIC-CXR, “SCOPE-Gen gen. only” achieves a mean set size of $6.003$ for the sum non-conformity in comparison to $7.390$ when using SCOPE-Gen with both filters. However, “SCOPE-Gen gen. only” is consistently more expensive in terms of admissibility queries than the standard version SCOPE-Gen, by a considerable gap. This can be explained by the fact that the filters require admissibility checks in a different order than the generation step, during calibration (visualized in Fig. 2). For instance, the quality filter orders the different examples from the previous prediction set according to their estimated quality. Typically, outputs with high quality estimates are also more likely to be admissible (we recall that we do not need to keep sampling beyond the first admissible example § 4.2). This lies in contrast to the order of samples considered in the calibration of the generation parameter, where such an ordering is not performed.
@@ E.2 Further Quantitative Evaluations
SCOPE-Gen flipped # Queries $4.760\pm 1.160$ $3.670\pm 0.570$-
Time $0.029\pm 0.013$ $0.025\pm 0.006$-
Set Size $7.744\pm 2.025$ $3.945\pm 0.995$-
@@
Appendix F Qualitative Results

### v2

@@ 7 Conclusion
Reproducibility Statement
Our source code is available on GitHub at https://github.com/rudolfwilliam/scope-gen. Instructions for reproducibility are provided in the README.md file at the top level of the code directory. All parameters for running the experiments can be found in the config folders within the script folders of each experiment folder.
Acknowledgements
Michael Muehlebach is funded by the German Research Foundation (DFG) under the project 456587626 of the Emmy Noether Programme. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Klaus-Rudolf Kladny.
@@ Appendix A Relationship between Risk Control and Admissibility Control
Appendix B Definition of the Empirical Quantile
We define the empirical $\delta$-quantile for $\delta\in\mathbb{R}$ at 1-dimensional real-valued data points $\{x_{i}\}_{i=1}^{n}$ as
$\widehat{\text{quantile}}\left(\{x_{i}\}_{i=1}^{n};\;\delta\right)\;\coloneqq\;\inf_{q\,\in\,\mathbb{R}\cup\{\infty\}}\left\{q\;:\;\sum_{i=1}^{n}\mathbbm{1}\{x_{i}\leq q\}\geq n\delta\right\}.$ (15)
Intuitively, 15 corresponds to the $n\delta^{\text{th}}$ smallest element in $\{x_{i}\}_{i=1}^{n}$ if $\delta\leq 1$. Otherwise ($\delta>1$), the empirical quantile 15 is set to $\infty$.
Appendix C Bounding Conformal Admissibility by Learn-then-Test Admissibility
@@
We note that 18 is bounded from below by 00. We further bound 17 as follows
@@
where the last inequality holds by the learn-then-test guarantee 16. ∎
Appendix D Implementation Details
D.1 Parameter Configuration
@@
where $K$ is the amount of prediction steps (typically, $K=3$). We note that 19 is somewhat analagous to a Bonferroni correction in that it also equally distributes a correction. In practice, when working with a upper bound on the sample size max (outlined in § 4.2), however, lower admissibility levels can often be attained by different choice of $\alpha_{(s)}$. Specifically, we choose a larger risk level for the generation step ($s=0$) and distribute the remaining risk equally among the other step(s) ($s\in\{1,2\}$). We assess two different configurations of SCOPE-Gen, which we name config. 1 and config 2. For config. 1, we choose $M=5$ and set
@@
For config. 2 , we use the same setup as for config. 1, with the only difference that we only use the diversity filter (i. e., $K=2$). In addition, we add a prediction set count penalty to the diversity update function, resulting in
$\texttt{update}(\nu,y)=\begin{cases}-d_{\text{max}},\quad&\text{if}\quad\nu=0\\ -\displaystyle\min_{y^{\prime}\in\mathcal{C}_{(s)}\setminus\{y\}}d(y,y^{\prime})+\gamma_{d}|\mathcal{C}_{(s)}|,\quad&\text{else},\end{cases}$
with $\gamma_{d}=0.1$.
D.2 Distance Upper Bound
@@ D.3 Grid Specification for CLM
Appendix E Experimental Details
E.1 Setup Details
@@
TriviaQA. We remove duplicates automatically rather than calibrating a parameter to do this, in order to improve statistical efficiency of both methods. For CLM, we set the diversity parameter $\lambda_{(2)}=0.5$ such that all duplicates are removed and for SCOPE-Gen, we use a fixed diversity filter at the end of the prediction pipeline, which removes duplicates.
@@ E.2 Examples
Appendix F Additional Experiments
F.1 Admissibility Analysis
We demonstrate the empirical admissibility levels for the MIMIC-CXR experiment at admissibility level $\alpha=0.3$, varying calibration set size $n$ and the used non-conformity measure. We use a cross-validation like algorithm (Angelopoulos & Bates (2021); Fig. 12) to generate histograms. The rest of the experimental setup is identical to the setup used for all other experiments (see caption of Tab. 1)
As shown in the histograms of Fig. 3, for all non-conformity measures, the admissibility achieved by SCOPE-Gen (config. 1 [§ D.1]; blue) is slightly conservative for low sample sizes and converges to the desired level as the calibration set size $n$ is increased from $300$ to $1200$. This is similar for CLM (orange), where the achieved admissibility (using the optimization strategy described in § 5) also becomes less conservative for large sample sizes. However, in contrast to SCOPE-Gen, CLM is generally more conservative. In addition, this conservativeness is highly influenced by the choice of non-conformity measure: While the CLM gap becomes tighter for the max and sum non-conformities, this is not the case for the count non-conformity, where a large gap persists even for $n=1200$.
@@ F.1 Admissibility Analysis
Figure 3: Admissibility Analysis for MIMIC-CXR. For all non-conformity scores, SCOPE-Gen (config. 1, [see § D.1]; blue) becomes less conservative with respect to the desired admissibility level $1-\alpha$ (red dashed line) as the amount of calibration samples $n$ increases, as typical for methods based on conformal prediction. For CLM (orange), in contrast, conservativeness depends much on the chosen non-conformity measure.
F.2 Further Quantitative Evaluations & Ablations
We demonstrate further experiments that showcase the performance for the different non-conformity measures (count and max ($\gamma=0.1$) in addition to sum ($\gamma=0.5$); see 10), including two more baselines (baselines 3) and 4) introduced in § 5) for admissibility levels $\alpha\in\{0.3,0.35,0.4\}$ and both configurations of SCOPE-Gen (see § D.1). While for some experiments, choosing smaller values for $\alpha$ is feasible (e.g., on TriviaQA, as can be seen in § G.1), this is not possible for all experiments due to the performance of the underlying model. Specifically, if a model does not generate a single admissible example within the upper limit max for more than $\alpha$ fraction of examples, we see that generating prediction sets that fulfill an admissibility of level smaller than $\alpha$ cannot be possible, irrespective of the calibration method that is used. We observe such issues for CNN/DM and Molecules when using admissibility levels smaller than $\alpha=0.3$. Thus, we restrict our numerical experiments to larger admissibility levels.
We see in LABEL:tab:performance_further that typically either SCOPE-Gen (config. 1 and config. 2, see § D.1) or one of its ablations performs best in terms of the considered evaluation metrics. We see that occasionally, SCOPE-Gen gen. only generates the smallest prediction sets on average. We hypothesize that this may have to do with the fact that SCOPE-Gen gen. will tend to be slightly less affected by variance in the estimation of the threshold, since SCOPE-Gen gen. has more data to estimate the latter. In terms of the non-conformity update function, we see that a general recommendation in favor of one over any other cannot be made. For example, considering the setting $\alpha=0.3$ for TriviaQA, we see that SCOPE-Gen config. 1 generates the smallest prediction sets for the sum non-conformity, but the fewest admissibility checks are achieved for the same method when using the count non-conformity. For $\alpha=0.3$ and MIMIX-CXR, SCOPE-Gen config. 2 generates the smallest prediction sets, using max non-conformity.
@@ F.2 Further Quantitative Evaluations & Ablations
SCOPE-Gen config. 2 # Queries -- $5.216\pm 0.710$ $4.128\pm 0.452$-
Time -- $0.023\pm 0.007$ $0.017\pm 0.003$-
Set Size -- $\bf{5.839}\pm 1.144$ $\bf{2.973}\pm 0.614$-
Frac. Reject -- $0.090\pm 0.000$ $0.017\pm 0.000$-
Set Size $2.050\pm 0.184$ $6.003\pm 1.040$ $3.302\pm 0.579$ $10.038\pm 0.984$
@@
SCOPE-Gen config. 2 # Queries -- $5.385\pm 0.732$ $4.162\pm 0.458$-
Time -- $0.015\pm 0.007$ $0.010\pm 0.003$-
Set Size -- $\bf{5.899}\pm 1.044$ $\bf{2.970}\pm 0.551$-
Frac. Reject -- $0.080\pm 0.000$ $0.013\pm 0.000$-
@@
SCOPE-Gen flipped # Queries -- $4.760\pm 1.160$ $3.670\pm 0.570$-
Time -- $0.029\pm 0.013$ $0.025\pm 0.006$-
Set Size -- $7.744\pm 2.025$ $3.945\pm 0.995$-
@@
SCOPE-Gen config. 2 # Queries -- $5.096\pm 0.744$ $4.199\pm 0.422$-
Time -- $0.021\pm 0.007$ $0.017\pm 0.002$-
Set Size -- $\bf{5.761}\pm 0.980$ $3.078\pm 0.575$-
Frac. Reject -- $0.160\pm 0.000$ $0.007\pm 0.000$-
@@
$\alpha=0.35$
Non-Conformity: sum
CLM # Queries $6.581\pm 0.210$ $20.000\pm 0.000$ $20.000\pm 0.000$ $19.269\pm 0.222$
Time $2.755\pm 0.028$ $106.192\pm 14.696$ $102.641\pm 12.754$ $15.574\pm 1.669$
Set Size $1.674\pm 0.224$ $11.240\pm 2.075$ $4.094\pm 1.118$ $13.573\pm 2.615$
Frac. Reject $\bf{0.000}\pm 0.000$ $0.017\pm 0.000$ $\bf{0.000}\pm 0.000$ $\bf{0.000}\pm 0.000$
CLM reduced max # Queries $4.212\pm 0.114$ $10.000\pm 0.000$ $10.000\pm 0.000$ $\bf{7.107}\pm 0.064$
Time $0.977\pm 0.009$ $23.966\pm 0.578$ $26.360\pm 1.630$ $0.874\pm 0.012$
Set Size $1.633\pm 0.443$ $13.774\pm 2.340$ $4.245\pm 1.165$ $9.280\pm 0.000$
Frac. Reject $\bf{0.000}\pm 0.000$ $0.187\pm 0.000$ $\bf{0.000}\pm 0.000$ $0.997\pm 0.000$
SCOPE-Gen config. 1 # Queries $\bf{2.757}\pm 0.151$ $\bf{3.881}\pm 0.213$ $3.352\pm 0.299$ $7.973\pm 0.536$
Time $0.026\pm 0.001$ $0.058\pm 0.005$ $0.055\pm 0.005$ $0.094\pm 0.003$
Set Size $\bf{1.072}\pm 0.069$ $4.345\pm 0.802$ $2.904\pm 0.553$ $7.468\pm 0.432$
SCOPE-Gen config. 2 # Queries -- $4.929\pm 0.436$ $3.835\pm 0.332$-
Time -- $0.019\pm 0.002$ $0.016\pm 0.001$-
Set Size -- $4.190\pm 0.707$ $\bf{2.083}\pm 0.354$-
Frac. Reject -- $0.003\pm 0.000$ $\bf{0.000}\pm 0.000$-
SCOPE-Gen gen. only # Queries $3.633\pm 0.036$ $6.518\pm 0.310$ $5.525\pm 0.297$ $9.641\pm 0.462$
Time $\bf{0.006}\pm 0.000$ $\bf{0.002}\pm 0.000$ $\bf{0.002}\pm 0.000$ $\bf{0.006}\pm 0.001$
Set Size $1.685\pm 0.010$ $\bf{3.979}\pm 0.557$ $2.163\pm 0.315$ $\bf{7.050}\pm 0.436$
SCOPE-Gen flipped # Queries -- $3.903\pm 0.119$ $\bf{2.994}\pm 0.102$-
Time -- $0.043\pm 0.002$ $0.041\pm 0.002$-
Set Size -- $4.475\pm 0.512$ $2.326\pm 0.140$-
Frac. Reject -- $\bf{0.000}\pm 0.000$ $\bf{0.000}\pm 0.000$-
Non-Conformity: count
CLM # Queries $6.562\pm 0.197$ $20.000\pm 0.000$ $20.000\pm 0.000$ $19.258\pm 0.218$
Time $2.300\pm 0.024$ $84.599\pm 9.140$ $82.832\pm 10.867$ $16.173\pm 1.491$
Set Size $3.855\pm 1.104$ $18.646\pm 1.637$ $15.203\pm 2.297$ $20.236\pm 3.853$
Frac. Reject $\bf{0.000}\pm 0.000$ $0.003\pm 0.000$ $\bf{0.000}\pm 0.000$ $\bf{0.000}\pm 0.000$
CLM reduced max # Queries $4.204\pm 0.114$ $10.000\pm 0.000$ $10.000\pm 0.000$ $\bf{7.105}\pm 0.070$
Time $0.469\pm 0.007$ $9.922\pm 0.088$ $10.170\pm 0.426$ $0.405\pm 0.008$
Set Size $3.819\pm 1.029$ $18.257\pm 1.709$ $13.249\pm 3.114$ $17.753\pm 0.000$
Frac. Reject $\bf{0.000}\pm 0.000$ $0.163\pm 0.000$ $\bf{0.000}\pm 0.000$ $0.997\pm 0.000$
SCOPE-Gen config. 1 # Queries $\bf{2.754}\pm 0.204$ $\bf{3.974}\pm 0.063$ $\bf{3.172}\pm 0.186$ $7.896\pm 0.485$
Time $0.011\pm 0.001$ $0.040\pm 0.002$ $0.033\pm 0.003$ $0.011\pm 0.001$
Set Size $\bf{1.055}\pm 0.128$ $4.569\pm 0.274$ $2.538\pm 0.451$ $7.723\pm 0.479$
SCOPE-Gen config. 2 # Queries -- $5.039\pm 0.449$ $3.924\pm 0.364$-
Time -- $0.011\pm 0.003$ $0.008\pm 0.001$-
Set Size -- $4.288\pm 0.703$ $\bf{2.226}\pm 0.371$-
Frac. Reject -- $\bf{0.000}\pm 0.000$ $\bf{0.000}\pm 0.000$-
SCOPE-Gen gen. only # Queries $3.633\pm 0.036$ $6.138\pm 0.119$ $5.525\pm 0.297$ $9.641\pm 0.462$
Time $\bf{0.005}\pm 0.000$ $\bf{0.004}\pm 0.000$ $\bf{0.002}\pm 0.000$ $\bf{0.005}\pm 0.000$
Set Size $1.722\pm 0.191$ $\bf{4.000}\pm 0.000$ $2.707\pm 0.491$ $\bf{7.074}\pm 0.486$
SCOPE-Gen flipped # Queries -- $4.750\pm 0.530$ $3.240\pm 0.240$-
Time -- $0.028\pm 0.004$ $0.021\pm 0.002$-
Set Size -- $5.815\pm 1.368$ $\bf{2.310}\pm 0.325$-
Frac. Reject -- $\bf{0.000}\pm 0.000$ $\bf{0.000}\pm 0.000$-
Non-Conformity: max
CLM # Queries $6.544\pm 0.203$ $20.000\pm 0.000$ $20.000\pm 0.000$ $19.267\pm 0.212$
Time $2.785\pm 0.021$ $111.025\pm 15.329$ $102.087\pm 12.566$ $15.301\pm 1.092$
Set Size $1.411\pm 0.190$ $10.699\pm 1.959$ $\bf{2.121}\pm 0.301$ $\bf{2.004}\pm 0.144$
CLM reduced max # Queries $4.206\pm 0.105$ $10.000\pm 0.000$ $10.000\pm 0.000$ $\bf{7.107}\pm 0.066$
Time $0.985\pm 0.010$ $24.432\pm 0.677$ $26.455\pm 1.557$ $0.872\pm 0.012$
Set Size $1.499\pm 0.304$ $13.781\pm 2.486$ $2.401\pm 0.426$ $2.420\pm 0.000$
Frac. Reject $\bf{0.000}\pm 0.000$ $0.150\pm 0.000$ $0.003\pm 0.000$ $0.997\pm 0.000$
SCOPE-Gen config. 1 # Queries $\bf{3.056}\pm 0.098$ $\bf{3.922}\pm 0.305$ $\bf{3.379}\pm 0.156$ $7.891\pm 0.532$
Time $0.025\pm 0.001$ $0.061\pm 0.008$ $0.055\pm 0.005$ $0.077\pm 0.003$
Set Size $\bf{1.334}\pm 0.076$ $4.317\pm 1.161$ $2.697\pm 0.144$ $7.598\pm 0.640$
SCOPE-Gen config. 2 # Queries -- $5.096\pm 0.744$ $4.199\pm 0.422$-
Time -- $0.021\pm 0.007$ $0.017\pm 0.002$-
Set Size -- $\bf{5.761}\pm 0.980$ $3.078\pm 0.575$-
Frac. Reject -- $0.160\pm 0.000$ $0.007\pm 0.000$-
SCOPE-Gen gen. only # Queries $3.633\pm 0.036$ $6.138\pm 0.119$ $5.525\pm 0.297$ $9.637\pm 0.462$
Time $\bf{0.006}\pm 0.000$ $\bf{0.005}\pm 0.000$ $\bf{0.002}\pm 0.000$ $\bf{0.006}\pm 0.001$
Set Size $3.024\pm 0.080$ $\bf{3.609}\pm 0.173$ $2.865\pm 0.307$ $7.025\pm 0.418$
SCOPE-Gen flipped # Queries -- $4.156\pm 0.129$ $3.384\pm 0.123$-
Time -- $0.043\pm 0.002$ $0.040\pm 0.002$-
Set Size -- $5.017\pm 0.743$ $3.315\pm 0.368$-
Frac. Reject -- $\bf{0.000}\pm 0.000$ $\bf{0.000}\pm 0.000$-
$\alpha=0.4$
Non-Conformity: sum
CLM # Queries $6.555\pm 0.196$ $20.000\pm 0.000$ $20.000\pm 0.000$ $19.247\pm 0.215$
Time $2.724\pm 0.018$ $107.511\pm 15.823$ $100.752\pm 12.215$ $15.054\pm 1.231$
Set Size $1.397\pm 0.258$ $8.105\pm 1.708$ $2.816\pm 0.500$ $9.952\pm 2.174$
CLM reduced max # Queries $4.207\pm 0.103$ $10.000\pm 0.000$ $10.000\pm 0.000$ $\bf{7.118}\pm 0.068$
Time $0.977\pm 0.009$ $23.955\pm 0.596$ $26.075\pm 1.545$ $0.864\pm 0.011$
Set Size $1.200\pm 0.219$ $10.039\pm 2.115$ $2.673\pm 0.550$ $9.639\pm 1.911$
Frac. Reject $\bf{0.000}\pm 0.000$ $0.003\pm 0.000$ $\bf{0.000}\pm 0.000$ $0.890\pm 0.000$
SCOPE-Gen config. 1 # Queries $\bf{2.674}\pm 0.147$ $\bf{3.724}\pm 0.343$ $\bf{2.698}\pm 0.033$ $7.600\pm 0.443$
Time $0.026\pm 0.001$ $0.052\pm 0.004$ $0.044\pm 0.002$ $0.096\pm 0.004$
Set Size $\bf{0.917}\pm 0.053$ $3.609\pm 0.666$ $\bf{1.379}\pm 0.058$ $5.864\pm 0.219$
SCOPE-Gen config. 2 # Queries $4.599\pm 0.355$ $3.653\pm 0.316$
Time $0.017\pm 0.001$ $0.015\pm 0.001$
Set Size $3.110\pm 0.454$ $1.578\pm 0.277$
Frac. Reject $\bf{0.000}\pm 0.000$ $\bf{0.000}\pm 0.000$
SCOPE-Gen gen. only # Queries $3.633\pm 0.036$ $6.518\pm 0.310$ $5.525\pm 0.297$ $9.646\pm 0.464$
Time $\bf{0.006}\pm 0.000$ $\bf{0.002}\pm 0.000$ $\bf{0.002}\pm 0.000$ $\bf{0.006}\pm 0.001$
Set Size $1.484\pm 0.021$ $\bf{2.889}\pm 0.348$ $1.586\pm 0.158$ $\bf{5.540}\pm 0.168$
SCOPE-Gen flipped # Queries -- $3.785\pm 0.156$ $2.775\pm 0.113$-
Time -- $0.042\pm 0.002$ $0.040\pm 0.003$-
Set Size -- $3.244\pm 0.468$ $1.486\pm 0.255$-
Frac. Reject -- $\bf{0.000}\pm 0.000$ $\bf{0.000}\pm 0.000$-
Non-Conformity: count
CLM # Queries $6.552\pm 0.193$ $20.000\pm 0.000$ $20.000\pm 0.000$ $19.261\pm 0.223$
Time $2.301\pm 0.026$ $87.296\pm 15.006$ $83.194\pm 10.589$ $15.572\pm 1.655$
Set Size $4.690\pm 1.879$ $19.155\pm 1.166$ $12.349\pm 1.964$ $16.278\pm 4.197$
CLM reduced max # Queries $4.200\pm 0.114$ $10.000\pm 0.000$ $10.000\pm 0.000$ $\bf{7.106}\pm 0.065$
Time $0.471\pm 0.007$ $10.006\pm 0.127$ $10.306\pm 0.534$ $0.409\pm 0.008$
Set Size $4.790\pm 1.840$ $19.281\pm 1.183$ $11.934\pm 2.491$ $14.779\pm 3.371$
Frac. Reject $\bf{0.000}\pm 0.000$ $\bf{0.000}\pm 0.000$ $\bf{0.000}\pm 0.000$ $0.833\pm 0.000$
SCOPE-Gen # Queries $\bf{2.661}\pm 0.099$ $\bf{3.667}\pm 0.076$ $\bf{2.937}\pm 0.197$ $7.509\pm 0.474$
Time $0.010\pm 0.000$ $0.032\pm 0.003$ $0.026\pm 0.003$ $0.011\pm 0.001$
Set Size $\bf{0.938}\pm 0.073$ $3.425\pm 0.449$ $\bf{1.708}\pm 0.257$ $6.009\pm 0.450$
SCOPE-Gen gen. only # Queries $3.633\pm 0.036$ $6.138\pm 0.119$ $5.525\pm 0.297$ $9.641\pm 0.462$
Time $\bf{0.005}\pm 0.000$ $\bf{0.004}\pm 0.000$ $\bf{0.002}\pm 0.000$ $\bf{0.005}\pm 0.001$
Set Size $1.553\pm 0.013$ $\bf{3.000}\pm 0.000$ $2.007\pm 0.081$ $\bf{5.625}\pm 0.272$
SCOPE-Gen flipped # Queries -- $3.910\pm 0.061$ $3.091\pm 0.475$-
Time -- $0.021\pm 0.001$ $0.019\pm 0.003$-
Set Size -- $3.495\pm 0.120$ $1.853\pm 0.414$-
Frac. Reject -- $\bf{0.000}\pm 0.000$ $\bf{0.000}\pm 0.000$-
Non-Conformity: max
CLM # Queries $6.574\pm 0.197$ $20.000\pm 0.000$ $20.000\pm 0.000$ $19.255\pm 0.210$
Time $2.816\pm 0.032$ $105.393\pm 12.837$ $101.326\pm 11.619$ $15.101\pm 0.937$
Set Size $\bf{1.170}\pm 0.155$ $7.871\pm 1.612$ $1.772\pm 0.209$ $\bf{1.927}\pm 0.097$
CLM reduced max # Queries $4.207\pm 0.120$ $10.000\pm 0.000$ $10.000\pm 0.000$ $\bf{7.111}\pm 0.066$
Time $0.987\pm 0.010$ $24.634\pm 0.634$ $26.286\pm 1.712$ $0.880\pm 0.013$
Set Size $1.176\pm 0.150$ $9.871\pm 2.349$ $1.838\pm 0.221$ $2.260\pm 0.219$
Frac. Reject $\bf{0.000}\pm 0.000$ $0.003\pm 0.000$ $\bf{0.000}\pm 0.000$ $0.857\pm 0.000$
SCOPE-Gen config. 1 # Queries $\bf{3.024}\pm 0.103$ $\bf{3.711}\pm 0.192$ $3.032\pm 0.089$ $7.515\pm 0.448$
Time $0.024\pm 0.001$ $0.052\pm 0.003$ $0.046\pm 0.002$ $0.076\pm 0.002$
Set Size $1.244\pm 0.098$ $3.622\pm 0.421$ $2.043\pm 0.078$ $5.942\pm 0.300$
SCOPE-Gen config. 2 # Queries -- $4.732\pm 0.398$ $3.703\pm 0.293$-
Time -- $0.008\pm 0.001$ $0.006\pm 0.001$-
Set Size -- $3.279\pm 0.520$ $\bf{1.708}\pm 0.263$-
Frac. Reject -- $\bf{0.000}\pm 0.000$ $\bf{0.000}\pm 0.000$-
SCOPE-Gen gen. only # Queries $3.633\pm 0.036$ $6.138\pm 0.119$ $5.525\pm 0.297$ $9.646\pm 0.464$
Time $\bf{0.005}\pm 0.000$ $\bf{0.005}\pm 0.001$ $\bf{0.002}\pm 0.000$ $\bf{0.006}\pm 0.002$
Set Size $2.935\pm 0.107$ $\bf{2.652}\pm 0.158$ $2.247\pm 0.203$ $5.586\pm 0.229$
SCOPE-Gen flipped # Queries -- $3.924\pm 0.375$ $\bf{3.019}\pm 0.178$-
Time -- $0.040\pm 0.002$ $0.037\pm 0.002$-
Set Size -- $4.083\pm 0.791$ $1.855\pm 0.145$-
Frac. Reject -- $\bf{0.000}\pm 0.000$ $\bf{0.000}\pm 0.000$-
Appendix G Qualitative Results